# Learning Discrete Concepts in Latent Hierarchical Models

### v1

@@ 1 Introduction
In natural images, the degree/extent of certain attributes (e.g., position, lighting) is often presented in a continuous form and main concepts of practical concern are often discrete in nature (e.g., object classes and shapes). Moreover, these concepts are often statistically dependent, with the dependence potentially resulting from some higher-level concepts. For example, the correlation of a specific dog’s eye features and fur features may arise from a high-level concept for breeds (Figure 1). Similarly, even higher-level concepts may exist and induce dependence between high-level concepts, giving rise to a hierarchical model that characterizes all discrete concepts at different abstraction levels underlying high-dimensional data distributions. In this work, we focus on concepts that can be defined as discrete latent variables and related via a hierarchical model. Under this formalization, the query on the recoverability of concepts and their relations from unstructured high-dimensional distribution (e.g., images) amounts to the following causal identification problem:
@@
Identification theory for latent hierarchical causal models has been a topic of sustained interest. Recent work [18, 19, 20] investigates identification conditions of latent hierarchical structures under the assumption that the latent variables are continuous and influence each other through linear functions. The linearity assumption fails to handle the general nonlinear influences among discrete variables. Another line of work focuses on discrete latent models. Pearl [21], Choi et al. [22] study latent trees with discrete observed variables. The tree structure can be over-simplified to capture the complex interactions among concepts from distinct abstract levels (e.g., multiple high-level concepts can jointly influence a lower-level one). Gu and Dunson [23] assume that binary latent variables can be exactly grouped into levels and causal edges often appear between adjacent levels, which can also be restrictive. Moreover, these papers assume observed variables are discrete, falling short of modeling the continuous distribution like images as the observed variables. Similar to our goal, Kivva et al. [24] show the discrete latent variables adjacent to the potentially continuous observed variables can be identified. However, their theory assumes the absence of higher-level latent variables and thus cannot handle latent hierarchical structures. More related work can be found in Appendix A1.
@@
As a implication of our theorems, we discuss a novel interpretation of the state-of-the-art latent diffusion (LD) models [28] through the lens of a hierarchical concept model. We interpret the denoising objective at different noise levels as estimating latent concept embeddings at corresponding hierarchical levels in the causal model, where a higher noise level corresponds to high-level concepts. This perspective explains and unifies these seemingly orthogonal threads of empirical insights and gives rise to insights for potential empirical improvements. We deduce several insights from our theoretical results and verify them empirically. In summary, our main contributions are as follows.
@@ 2 Discrete Hierarchical Models
Figure 1: Latent causal models. The dashed circle in (a) indicates that the continuous variable $\mathbf{c}$ can be viewed as an exogenous variable. Dashed edges in (b) indicate potential statistical dependence.
@@
Given this, we define the discrete hierarchical model as follows. The discrete hierarchical model (Figure 1(c)) $\mathcal{G}:=(\mathbf{E},\mathbf{V})$ is a DAG that comprises discrete latent variables $d_{1},\cdots,d_{n_{d}},z_{1},\cdots,z_{n_{z}}$. We denote that directed edge set with $\mathbf{E}$ and the collection of all variables with $\mathbf{V}:=\{\mathbf{D},\mathbf{Z}\}$, where $\mathbf{D}$ and $\mathbf{Z}$ are vectors $\mathbf{d}$ and $\mathbf{z}$ in a set form and all leaf variables in $\mathcal{G}$ belong to $\mathbf{D}$. We assume the distribution over all variables $\mathbf{V}$ respects the Markov property with respect to the graph $\mathcal{G}$. We denote all parents and children of a variable with $\text{Ch}(\cdot)$ and $\text{Pa}(\cdot)$ respectively. If a set of variables $\mathbf{A}\subset\mathbf{V}$ has no other parents than $\mathbf{B}$, we say $\mathbf{A}$ is the pure children of $\mathbf{B}$ (Definition A8), i.e., $\mathbf{A}=\text{PCh}_{\mathcal{G}}(\mathbf{B})$. As shown in Figure 1(c), $x_{1}$ is a pure child of $d_{1}$.
@@ 3 Identification of Discrete Latent Hierarchical Models
We denote the set containing components of $\mathbf{x}$ with $\mathbf{X}$, the set of all variables with $\mathbf{V}^{*}:=\mathbf{V}\cup\mathbf{X}$, the entire edge set with $\mathbf{E}^{*}:=\mathbf{E}\cup\Gamma$, and the entire causal model with $\mathcal{G}^{*}:=(\mathbf{V}^{*},\mathbf{E}^{*})$. As the true generating process involves $\mathbf{d}$, $\mathbf{c}$, $g$, $\Gamma$, and $\mathcal{G}$ (defined in Section 2), we define their statistical estimates with $\hat{\mathbf{d}}$, $\hat{\mathbf{c}}$, $\hat{g}$, and $\hat{\Gamma}$ through maximum likelihood estimation over the observed distribution $p(\mathbf{x})$ while respecting conditions on the true generating process.
@@ 3.1 General Conditions for Discrete Latent Models
Condition 3.1 (General Latent Model Conditions).
@@
    [No-twins]: Distinct latent variables have distinct neighbors $\text{ne}(V_{1})\neq\text{ne}(v_{2})$, if $v_{i}\neq v_{j}\in\mathbf{V}$.
@@
    [Maximality]: There is no DAG $\tilde{\mathcal{G}}^{*}:=(\tilde{\mathbf{V}}^{*},\tilde{\mathbf{E}}^{*})$ resulting from splitting a latent variable in $\mathcal{G}^{*}$, such that ${\mathbb{P}}\left(\tilde{\mathbf{V}}^{*}\right)$ is Markov w.r.t. $\tilde{\mathcal{G}}^{*}$ and $\tilde{\mathcal{G}}^{*}$ satisfies ii.
@@ Discussion.
Condition 3.1 is a necessary set of conditions for identifying latent discrete models, which is employed and discussed extensively [24, 29]. Intuitively, Condition 3.1-i excludes dummy discrete states and graph edges that exert no influence on the observed variables $\mathbf{x}$. Condition 3.1-ii,iii constrain the latent model to be the most informative graph without introducing redundant latent variables, thus forbidding arbitrary merging and splitting over latent variables.
@@ 3.2 Discrete Component Identification
Condition 3.3 (Discrete Components Identification).
@@
    [Open & Connected Spaces] The continuous support $\mathcal{C}\subset\mathbb{R}^{n_{c}}$ is open and connected.
@@
Condition 3.3-i requires the continuous support $\mathcal{C}$ to be regular in contrast with the discrete variable’s support. Intuitively, the continuous variable $\mathbf{c}$ characterize all the smooth variations in the data, leaving the discontinuities explained by $\mathbf{d}$. Condition 3.3-ii ensures the generating process preserves latent variables’ information, without which the recovery of latent variables $\mathbf{d}$ is ill-posed. This condition is widely adopted in latent variable identification literature [30, 33, 34, 35]. Condition 3.3-iii ensures that each latent component should exhibit sufficiently distinguishable influences on the observed variable $\mathbf{x}$. It is also adopted in Kivva et al. [24, 29] and considered much weaker than assuming the existence of pure children for each latent variable [36, 37].
@@
Under the generating process in Equation 1 and Condition 3.3-ii, the estimated discrete variable $\hat{\mathbf{d}}$ and the true discrete variable $\mathbf{d}$ are equivalent up to an invertible function, i.e., $\hat{\mathbf{d}}=h(\mathbf{d})$ with $h(\cdot)$ invertible. Moreover, if Condition 3.1 and Condition 3.3-iii further hold, we attain component-wise identifiability (Definition 3.2) and the bipartite graph $\Gamma$ up to permutation of component indices.
@@
Intuitively, each state of the discrete subspace $\mathbf{d}$ indexes a manifold $g(\mathbf{d},\cdot):\mathbf{c}\mapsto\mathbf{x}$ that maps the continuous subspace $\mathbf{c}$ to the observed variable $\mathbf{x}$. These manifolds do not intersect in the observed variable space $\mathcal{X}$ regardless of however close they may be to each other, thanks to the invertibility of the generating function $g$ (Condition 3.3-ii). This leaves a sufficient footprint in $\mathbf{x}$ for us to uniquely identify the manifold it resides in, giving rise to the identifiability of $\mathbf{d}$. This reveals the discrete state of each realization of $\mathbf{x}$ and equivalently the joint distribution $p(\tilde{d},\mathbf{x})$ where we merge all components in $\mathbf{d}$ into a one-dimensional discrete variable $\tilde{d}$. Identifying this joint distribution enables the application of tensor decomposition techniques [24] to disentangle the global state $\hat{d}$ into individual discrete components $d_{i}$ and the associated causal graph $\Gamma$, under Condition 3.1 and Condition 3.3-iii. A detailed proof is given in Appendix A2.
@@ 3.3 Hierarchical Model Identification
Benefiting from the identified discrete components in Theorem 3.4, we employ $\mathbf{d}$ as observed variables to identify the discrete latent hierarchical model $\mathcal{G}$. Although discrete latent hierarchical models have been under investigation for an extensive period, existing results mostly assume relatively strong graphical conditions – the causal structures are either trees [26, 21, 22] or multi-level DAGs [23, 38], which can be restrictive in capturing the complex interactions among latent variables among different hierarchical levels. Separately, recent work [19, 20] has exhibited more flexible graphical conditions for linear, continuous latent hierarchical models. For instance, Dong et al. [20] allow for multiple directed paths of disparate edge numbers within a variable pair and potential non-leaf observed variables. Unfortunately, their techniques hinge on linearity and cannot directly apply to discrete models of high nonlinearity.
@@
To present Theorem 3.5, we introduce non-negative ranks $\text{rank}_{+}(\cdot)$ [40], a form of matrix ranks with additional structures (see Definition A5). We use $\left\lvert\text{Supp}(\mathbf{L})\right\rvert$ for the cardinality of a discrete variable set $\mathbf{L}$’s support and $\mathbf{P}_{\mathbf{A},{\mathbf{B}}}$ for the joint probability table whose two dimensions are the states of discrete variable sets $\mathbf{A}$ and $\mathbf{B}$ respectively. We define t-separation in Definition A2.
@@
Given two sets of variables $\mathbf{A}$ and $\mathbf{B}$ from a non-degenerate, faithful (Condition 3.1-i, Condition 3.7-i) discrete model $\mathcal{G}$, it follows that $\text{rank}_{+}(\mathbf{P}_{\mathbf{A},{\mathbf{B}}})=\min\{\left\lvert\text{Supp}(\mathbf{L})\right\rvert:\mathbf{L}~{}\text{t-separates}~{}\mathbf{A}~{}\text{and}~{}\mathbf{B}~{}\text{in}~{}\mathcal{G}\}$.

### v2

@@ 1 Introduction
In natural images, the degree/extent of certain attributes (e.g., position, lighting) is often presented in a continuous form and main concepts of practical concern are often discrete in nature (e.g., object classes and shapes). Moreover, these concepts are often statistically dependent, with the dependence potentially resulting from some higher-level concepts. For example, the correlation between a specific dog’s eye features and fur features may arise from a high-level concept for breeds (Figure 1). Similarly, even higher-level concepts may exist and induce dependence between high-level concepts, giving rise to a hierarchical model that characterizes all discrete concepts at different abstraction levels underlying high-dimensional data distributions. In this work, we focus on concepts that can be defined as discrete latent variables and related via a hierarchical model. Under this formalization, the query on the recoverability of concepts and their relations from unstructured high-dimensional distribution (e.g., images) amounts to the following causal identification problem:
@@
Identification theory for latent hierarchical causal models has been a topic of sustained interest. Recent work [18, 19, 20] investigates identification conditions of latent hierarchical structures under the assumption that the latent variables are continuous and influence each other through linear functions. The linearity assumption fails to handle the general nonlinear influences among discrete variables. Another line of work focuses on discrete latent models. Pearl [21], Choi et al. [22] study latent trees with discrete observed variables. The tree structure can be over-simplified to capture the complex interactions among concepts from distinct abstract levels (e.g., multiple high-level concepts can jointly influence a lower-level one). Gu and Dunson [23] assume that binary latent variables can be exactly grouped into levels and causal edges often appear between adjacent levels, which can also be restrictive. Moreover, these papers assume observed variables are discrete, falling short of modeling the continuous distribution like images as the observed variables. Similar to our goal, Kivva et al. [24] show the discrete latent variables adjacent to the potentially continuous observed variables can be identified. However, their theory assumes the absence of higher-level latent variables and thus cannot handle latent hierarchical structures.
@@
As an implication of our theorems, we discuss a novel interpretation of the state-of-the-art latent diffusion (LD) models [28] through the lens of a hierarchical concept model. We interpret the denoising objective at different noise levels as estimating latent concept embeddings at corresponding hierarchical levels in the causal model, where a higher noise level corresponds to high-level concepts. This perspective explains and unifies these seemingly orthogonal threads of empirical insights and gives rise to insights for potential empirical improvements. We deduce several insights from our theoretical results and verify them empirically. In summary, our main contributions are as follows.
@@
2 Related Work
Concept learning.
In recent years, a significant strand of research has focused on employing labeled data to learn concepts in generative models’ latent space for image editing and manipulation [1, 2, 3, 4, 5, 6]. Concurrently, another independent research trajectory has been exploring unsupervised concept discovery and its potential to learn more compositional and transferable models [7, 8, 9, 10, 11]. Concurrently, a plethora of work has been dedicated to extracting interpretable concepts from high-dimensional data such as images. Concept-bottleneck [12] first predicts a set of human-annotated concepts as an intermediate stage and then predicts the task labels from these intermediate concepts. This paradigm has attracted a large amount of follow-up work [13, 29, 30, 31, 32, 33]. A recent surge of pre-trained multimodal models (e.g., CLIP [17]) can explain the image concepts through text directly [14, 15, 16].
Latent variable identification.
Complex real-world data distributions often possess a hierarchical structure among their underlying latent variables. The identification conditions of latent hierarchical structures are investigated under the assumption that the latent variables are continuous and influence each other through linear functions [18, 19, 20] and nonlinear functions [34]. In addition, prior work [21, 26, 22, 23] studies fully discrete cases and thus falls short of modeling the continuous observed variables like images. To identify latent variables under nonlinear transformations, a line of work [35, 36, 37, 38] assumes the availability of auxiliary information (e.g., domain/class labels) and that the latent variables’ probability density functions have sufficiently different derivatives over domains/classes. Another line of studies [39, 40] refrains from the auxiliary information by assuming sparsity and mechanistic independence, disregarding causal structures among the latent variables.
Please refer to Section A1 for more extensive related work and discussion.
@@ 3 Discrete Hierarchical Models
Figure 1: Latent hierarchical graphs. The dashed circle in (a) indicates that the continuous variable $\mathbf{c}$ can be viewed as an exogenous variable. Dashed edges in (b) indicate potential statistical dependence.
@@
Given this, we define the discrete hierarchical model as follows. The discrete hierarchical model (Figure 1(c)) $\mathcal{G}:=(\mathbf{E},\mathbf{V})$ is a DAG that comprises discrete latent variables $d_{1},\cdots,d_{n_{d}},z_{1},\cdots,z_{n_{z}}$. We denote that directed edge set with $\mathbf{E}$ and the collection of all variables with $\mathbf{V}:=\{\mathbf{D},\mathbf{Z}\}$, where $\mathbf{D}$ and $\mathbf{Z}$ are vectors $\mathbf{d}$ and $\mathbf{z}$ in a set form and all leaf variables in $\mathcal{G}$ belong to $\mathbf{D}$. We assume the distribution over all variables $\mathbf{V}$ respects the Markov property with respect to the graph $\mathcal{G}$. We denote all parents and children of a variable with $\text{Ch}(\cdot)$ and $\text{Pa}(\cdot)$ respectively and define the neighbors as $\text{ne}(\cdot):=\text{Ch}(\cdot)\cup\text{Pa}(\cdot)$. We say a variable set $\mathbf{A}$ are pure children of $\mathbf{B}$, iff $\text{Pa}_{\mathcal{G}}(\mathbf{A})=\cup_{A_{i}\in\mathbf{A}}\text{Pa}_{\mathcal{G}}(A_{i})=\mathbf{B}$ and $\mathbf{A}\cap\mathbf{B}=\emptyset$. As shown in Figure 1(c), $x_{1}$ is a pure child of $d_{1}$.
@@ 4 Identification of Discrete Latent Hierarchical Models
We denote the set containing components of $\mathbf{x}$ with $\mathbf{X}$, the set of all variables with $\mathbf{V}^{*}:=\mathbf{V}\cup\mathbf{X}$, the entire edge set with $\mathbf{E}^{*}:=\mathbf{E}\cup\Gamma$, and the entire causal model with $\mathcal{G}^{*}:=(\mathbf{V}^{*},\mathbf{E}^{*})$. As the true generating process involves $\mathbf{d}$, $\mathbf{c}$, $g$, $\Gamma$, and $\mathcal{G}$ (defined in Section 3), we define their statistical estimates with $\hat{\mathbf{d}}$, $\hat{\mathbf{c}}$, $\hat{g}$, and $\hat{\Gamma}$ through maximum likelihood estimation over the full population $p(\mathbf{x})$ while respecting conditions on the true generating process. We use $\left\lvert\text{Supp}(\mathbf{L})\right\rvert$ for the cardinality of a discrete variable set $\mathbf{L}$’s support (all joint states) and $\mathbf{P}_{\mathbf{A},{\mathbf{B}}}$ for the joint probability table whose two dimensions are the states of discrete variable sets $\mathbf{A}$ and $\mathbf{B}$ respectively.
@@ 4.1 General Conditions for Discrete Latent Models
Condition 4.1 (General Latent Model Conditions).
@@
    [No-twins]: Distinct latent variables have distinct neighbors $\text{ne}(v_{1})\neq\text{ne}(v_{2})$, if $v_{1}\neq v_{2}\in\mathbf{V}$.
@@
    [Maximality]: There is no DAG $\tilde{\mathcal{G}}^{*}:=(\tilde{\mathbf{V}}^{*},\tilde{\mathbf{E}}^{*})$ resulting from splitting a latent variable in $\mathcal{G}^{*}$ (i.e., turning $z_{i}$ into $\tilde{z}_{i,1}$ and $\tilde{z}_{i,2}$ with identical neighbors and cardinality $|\Omega^{z}_{i}|=|\tilde{\Omega}^{z}_{i,1}|+|\tilde{\Omega}^{z}_{i,2}|$ ), such that ${\mathbb{P}}\left(\tilde{\mathbf{V}}^{*}\right)$ is Markov w.r.t. $\tilde{\mathcal{G}}^{*}$ and $\tilde{\mathcal{G}}^{*}$ satisfies ii.
@@ Discussion.
Condition 4.1 is a necessary set of conditions for identifying latent discrete models, which is employed and discussed extensively [24, 41]. Intuitively, Condition 4.1-i excludes dummy discrete states and graph edges that exert no influence on the observed variables $\mathbf{x}$. Condition 4.1-ii,iii constrain the latent model to be the most informative graph without introducing redundant latent variables, thus forbidding arbitrary merging and splitting over latent variables.
@@ 4.2 Discrete Component Identification
Condition 4.3 (Discrete Components Identification).
@@
    [Connected Spaces] The continuous support $\mathcal{C}\subset\mathbb{R}^{n_{c}}$ is closed and connected.
@@
Condition 4.3-i requires the continuous support $\mathcal{C}$ to be regular in contrast with the discrete variable’s support. Intuitively, the continuous variable $\mathbf{c}$ often controls the extents/degrees of specific attributes (e.g., sizes, lighting, and angles) and takes values from connected spaces. For instance, “lightning” ranges from the lowest to the highest intensity continuously. Condition 4.3-ii ensures the generating process preserves latent variables’ information [37, 35, 42, 43]. Thanks to the high dimensionality, images often have adequate capacity to meet this condition. For instance, the image of a dog contains a detailed description of the dog’s breed, shape, color, lighting intensity, and angles, all of which are decodable from the image. Condition 4.3-iii ensures that each latent component should exhibit sufficiently distinguishable influences on the observed variable $\mathbf{x}$. Practically, this condition indicates that the lowest-level concepts influence diverse parts of the image. These concepts are often atomic, such as a dog’s ear, eyes, or even finer, which often don’t overlap. This condition is adopted in prior work [24, 41] and related to the notation of sparsity. Along this line, prior work [44, 45, 46] assumes pure observed children for each discrete variable, which is strictly stronger. Recent work [47] assumes each latent variable is connected to a unique set of observed variables. This condition implies Condition 4.3-iii because if $z_{0}$’s children form a subset of $z_{1}$’s children, then one cannot find a subset of observed variables whose parent is $z_{0}$ alone.
@@
Under the generating process in Equation 1 and Condition 4.3-ii, the estimated discrete variable $\hat{\mathbf{d}}$ and the true discrete variable $\mathbf{d}$ are equivalent up to an invertible function, i.e., $\hat{\mathbf{d}}=h(\mathbf{d})$ with $h(\cdot)$ invertible. Moreover, if Condition 4.1 and Condition 4.3-iii further hold, we attain component-wise identifiability (Definition 4.2) and the bipartite graph $\Gamma$ up to permutation of component indices.
@@
Intuitively, each state of the discrete subspace $\mathbf{d}$ indexes a manifold $g(\mathbf{d},\cdot):\mathbf{c}\mapsto\mathbf{x}$ that maps the continuous subspace $\mathbf{c}$ to the observed variable $\mathbf{x}$. These manifolds do not intersect in the observed variable space $\mathcal{X}$ regardless of however close they may be to each other, thanks to the invertibility of the generating function $g$ (Condition 4.3-ii). This leaves a sufficient footprint in $\mathbf{x}$ for us to uniquely identify the manifold it resides in, giving rise to the identifiability of $\mathbf{d}$. This reveals the discrete state of each realization of $\mathbf{x}$ and equivalently the joint distribution $p(\tilde{d},\mathbf{x})$ where we merge all components in $\mathbf{d}$ into a discrete variable $\tilde{d}$. Identifying this joint distribution enables the application of tensor decomposition techniques [24] to disentangle the global state $\hat{d}$ into individual discrete components $d_{i}$ and the causal graph $\Gamma$, under Condition 4.1 and Condition 4.3-iii.
@@ 4.3 Hierarchical Model Identification
Benefiting from the identified discrete components in Theorem 4.4, we employ $\mathbf{d}$ as observed variables to identify the discrete latent hierarchical model $\mathcal{G}$. Although discrete latent hierarchical models have been under investigation for an extensive period, existing results mostly assume relatively strong graphical conditions – the causal structures are either trees [26, 21, 22] or multi-level DAGs [23, 48], which can be restrictive in capturing the complex interactions among latent variables among different hierarchical levels. Separately, recent work [19, 20] has exhibited more flexible graphical conditions for linear, continuous latent hierarchical models. For instance, prior work [20] allows for multiple directed paths of disparate edge numbers within a variable pair and potential non-leaf observed variables. Unfortunately, their techniques hinge on linearity and cannot directly apply to discrete models of high nonlinearity.
@@
To present Theorem 4.8, we introduce non-negative rank $\text{rank}_{+}(\cdot)$ [50] (Definition 4.5), and t-separation [51] (Definition 4.7) as follows.
Definition 4.5 (Non-negative Rank).
The non-negative rank of non-negative $A\in\mathbb{R}_{+}^{m\times n}$ is equal to the smallest $p$ for which there exist $\mathbf{B}\in\mathbb{R}^{m\times p}_{+}$ and $\mathbf{C}\in\mathbb{R}_{+}^{p\times n}$ such that $\mathbf{A}=\mathbf{B}\mathbf{C}$.
Definition 4.6 (Treks).
A trek $T_{i,j}$ in a DAG from vertex $i$ to $j$ consists of a directed path $P_{ki}$ from $k$ to $i$ and a direct path $P_{kj}$ from $k$ to $j$, where we refer to $P_{ki}$ as the $i$ side and $P_{kj}$ as the $j$ side.
Definition 4.7 (T-separation).
Intuitively, a trek is a path containing at most one fork structure and no collider structures. It is known that one can formulate d-separation as a special form of t-separation (see Theorem A1). Thus, t-separation is at least as informative as d-separation. As detailed in Dong et al. [20], t-separation can provide more information when latent variables are involved, benefiting from Theorem A2 [49].
@@
Given two sets of variables $\mathbf{A}$ and $\mathbf{B}$ from a non-degenerate, faithful (Condition 4.1-i, Condition 4.10-i) discrete model $\mathcal{G}$, it follows that $\text{rank}_{+}(\mathbf{P}_{\mathbf{A},{\mathbf{B}}})=\min\{\left\lvert\text{Supp}(\mathbf{L})\right\rvert:\text{a partition }(\mathbf{L}_{1},\mathbf{L}_{2})\text{ t-separates}~{}\mathbf{A}~{}\text{and}~{}\mathbf{B}~{}\text{in}~{}\mathcal{G}\}$.